# GRAND: Graph Release with Assured Node Differential Privacy

### v2

@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1$
For any $(x_{1},x_{2})$, when $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we can bound: Note that $|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|\leq|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|$. Moreover, by moving the sup operator from outside the sum to inside the sum, we have:
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})}[|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}[\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\sup_{(x_{1},x_{2})}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$=\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}[\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:|\bar{Z}_{i1}-T_{i1}|<\varepsilon}\sup_{x_{1}}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}|+\frac{1}{n}\sum_{i:|\bar{Z}_{i2}-T_{i2}|<\varepsilon}\sup_{x_{2}}|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}|.$
For each coordinate $j=1,2$, the term $\sup_{x_{j}}|\mathbf{1}_{\{\bar{Z}_{ij}\leq x_{j}\}}-\mathbf{1}_{\{T_{ij}\leq x_{j}\}}|$ equals 1 only if there exists some $x_{j}$ between $\bar{Z}_{ij}$ and $T_{ij}$. When $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we have $|\bar{Z}_{ij}-T_{ij}|<\varepsilon$ for both $j=1,2$.
@@
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}|=O_{P,n}(\varepsilon).$
@@
$\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\iff|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}$
@@
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}|\hat{F}_{m}(Z_{i1})-F_{1}(Z_{i1})|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
@@
$\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}|\hat{F}_{m}^{-1}(u)-F_{1}^{-1}(u)|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
where $\delta,\delta^{\prime}$ are chosen as in Lemma C.1, and $[a,b]$ contains the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$.
For the second coordinate, conditioning on $\bar{Z}_{i1}$ and $T_{i1}$, define analogous events. Specifically, let $a,b$ be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $U(0,1)$. Define
$\mathscr{E}_{m,2}=\left\{\sup_{x_{1}}\sup_{1\leq i\leq n}|\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})-F_{2|1}(Z_{i2}|Z_{i1})|<\delta\right\}\cap\left\{\sup_{x_{1}}\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(v|x_{1})|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
From Lemma B.2 we know that $P(\mathscr{E}_{m,2})\to 1$. Similarly, define
$\mathscr{E}_{m,2}^{\prime}=\left\{\sup_{x_{1}}\sup_{u\in[a,b]}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-F_{2|1}^{-1}(u|x_{1})|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
@@
$\{\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\}$ $\subseteq\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\}$
$\subseteq\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\}\cup\{|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\}$
@@
$P\left(\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i1}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\to 0$
$P\left(\frac{1}{n}\sum_{i}\mathbf{1}_{|\{U_{i2}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\to 0$
@@
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i}[\mathbf{1}_{|\{U_{i1}\notin[a,b]\}}+\mathbf{1}_{|\{U_{i2}\notin[a,b]\}}]$
@@
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\tilde{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})|\xrightarrow{P}0$
@@
$\sup_{(x_{1},x_{2})\in{\bf R}^{2}}|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})|\xrightarrow{P}0.$
For any $(x_{1},x_{2})$:
$|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})|$ $=\Big{|}\frac{1}{n}\sum_{i=1}^{n}[\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}]\Big{|}$
@@
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}|$
$+\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1$
@@
$\sup_{(x_{1},x_{2})}\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}|$
$\leq\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}[\sup_{x_{1}}|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}|+\sup_{x_{2}}|\mathbf{1}_{\{\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}|]$
Using our assumptions about the true $F$ and because of the convergence indicated by Lemma C.3, this term is $O_{P,n}(\varepsilon)$.
@@
$\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon\iff|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}$
@@
$\frac{1}{n}\sum_{i}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}\leq\gamma$
@@
$|\tilde{Z}_{i2}-\bar{Z}_{i2}|$ $=|\hat{F}_{2|1,m}^{-1}(u_{i}|\tilde{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(v_{i}|\bar{Z}_{i1})|$
$\leq|\hat{F}_{2|1,m}^{-1}(u_{i}|\tilde{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(u_{i}|\bar{Z}_{i1})|+|\hat{F}_{2|1,m}^{-1}(u_{i}|\bar{Z}_{i1})-\hat{F}_{2|1,m}^{-1}(v_{i}|\bar{Z}_{i1})|$
@@
$P(\frac{1}{n}\sum_{i}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}>\gamma/4)\to 0,\quad P(\frac{1}{n}\sum_{i}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}>\gamma/4)\to 0.$
@@
$\sup_{x_{1}\in K}\sup_{\begin{subarray}{c}y_{1}\in K\\ |x_{1}-y_{1}|<\delta\end{subarray}}\sup_{u\in[a,b]}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(u|y_{1})|<\varepsilon/2\sqrt{2}$
@@
$\sup_{x_{1}\in K}\sup_{u,v\in[a,b],|u-v|<\delta_{1}}|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(v|x_{1})|<\varepsilon/2\sqrt{2}$
@@
$\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\}\subseteq\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\}\cup\{u_{i}\notin[a,b]\}\cup\{v_{i}\notin[a,b]\}\cup\{|u_{i}-v_{i}|\geq\delta_{1}\}$

### v3

@@ Appendix C Proof of Theorem 3 (Emprical CDF Convergence)
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\big|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1.$
For any $(x_{1},x_{2})\in\mathbb{R}^{2}$, when $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we can bound: Note that $\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\big|\leq\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|$. Moreover, by moving the sup operator from outside the sum to inside the sum, we have:
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\big|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\big|$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\left[\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|\right]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\left[\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|\right]$
$=\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\left[\sup_{x_{1}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\sup_{x_{2}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|\right]$
$\leq\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{1}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\sup_{x_{2}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|$
$\leq\frac{1}{n}\sum_{i:|\bar{Z}_{i1}-T_{i1}|<\varepsilon}\sup_{x_{1}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1}\}}\big|+\frac{1}{n}\sum_{i:|\bar{Z}_{i2}-T_{i2}|<\varepsilon}\sup_{x_{2}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i2}\leq x_{2}\}}\big|.$
For each coordinate $j=1,2$, the term $\sup_{x_{j}\in\mathbb{R}}\big|\mathbf{1}_{\{\bar{Z}_{ij}\leq x_{j}\}}-\mathbf{1}_{\{T_{ij}\leq x_{j}\}}\big|$ equals 1 only if there exists some $x_{j}$ between $\bar{Z}_{ij}$ and $T_{ij}$. When $\|\bar{Z}_{i}-T_{i}\|<\varepsilon$, we have $|\bar{Z}_{ij}-T_{ij}|<\varepsilon$ for both $j=1,2$.
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|<\varepsilon}\big|\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{T_{i1}\leq x_{1},T_{i2}\leq x_{2}\}}\big|=\mathcal{O}_{\mathbb{P},n}(\varepsilon).$
@@
$\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\implies|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}.$
@@
$\mathscr{E}_{m}=\left\{\sup_{1\leq i\leq n}\big|\hat{F}_{m}(Z_{i1})-F_{1}(Z_{i1})\big|<\delta\right\}\cap\left\{\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{m}^{-1}(u)-\hat{F}_{m}^{-1}(v)\big|<\frac{\varepsilon}{2\sqrt{2}}\right\}$
@@
$\mathscr{E}_{m}^{\prime}=\left\{\sup_{u\in[a,b]}\big|\hat{F}_{m}^{-1}(u)-F_{1}^{-1}(u)\big|<\frac{\varepsilon}{2\sqrt{2}}\right\}.$
where $\delta,\delta^{\prime}$ are chosen as in Lemma C.1, and $[a,b]$ contains the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $\text{Uniform}(0,1)$.
For the second coordinate, conditioning on $\bar{Z}_{i1}$ and $T_{i1}$, define analogous events. Specifically, let $a,b$ be the $\varepsilon/16$ and $1-\varepsilon/16$ quantiles of $\text{Uniform}(0,1)$. Define
$\mathscr{E}_{m,2}$ $=\left\{\sup_{1\leq i\leq n}\big|\hat{F}_{2|1,m}(Z_{i2}|Z_{i1})-F_{2|1}(Z_{i2}|Z_{i1})\big|<\delta\right\}$
$\quad\cap\left\{\sup_{x_{1}\in\mathbb{R}}\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta^{\prime}\end{subarray}}\big|\hat{F}_{2|1,m}^{-1}(u|x_{1})-\hat{F}_{2|1,m}^{-1}(v|x_{1})\big|<\frac{\varepsilon}{2\sqrt{2}}\right\}.$
From Lemma B.2 we know that $\mathbb{P}(\mathscr{E}_{m,2})\xrightarrow{m\to\infty}1$. Similarly, define
$\mathscr{E}_{m,2}^{\prime}=\left\{\sup_{x_{1}\in\mathbb{R}}\sup_{u\in[a,b]}\big|\hat{F}_{2\,|\,1,m}^{-1}(u\,|\,x_{1})-F_{2\,|\,1}^{-1}(u\,|\,x_{1})\big|<\frac{\varepsilon}{2\sqrt{2}}\right\}.$
@@
$\left\{\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon\right\}$ $\subseteq\left\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\right\}$
$\subseteq\left\{|\bar{Z}_{i1}-T_{i1}|\geq\varepsilon/\sqrt{2}\right\}\cup\left\{|\bar{Z}_{i2}-T_{i2}|\geq\varepsilon/\sqrt{2}\right\}.$
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{U_{i1}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\xrightarrow{n\to\infty}0$
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{U_{i2}\notin[a,b]\}}>\frac{\varepsilon}{4}\right)\xrightarrow{n\to\infty}0$
@@
$\frac{1}{n}\sum_{i:\|\bar{Z}_{i}-T_{i}\|\geq\varepsilon}1$ $\leq\frac{1}{n}\sum_{i=1}^{n}\left[\mathbf{1}_{\{U_{i1}\notin[a,b]\}}+\mathbf{1}_{\{U_{i2}\notin[a,b]\}}\right]$
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\tilde{F}_{n}(x_{1},x_{2})-F(x_{1},x_{2})\big|\xrightarrow{\mathbb{P}}0\quad\text{as }m,n\to\infty,$
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\big|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})\big|\xrightarrow{\mathbb{P}}0.$
For any $(x_{1},x_{2})\in\mathbb{R}^{2}$:
$\big|\tilde{F}_{n}(x_{1},x_{2})-\bar{F}_{n}(x_{1},x_{2})\big|$ $=\left|\frac{1}{n}\sum_{i=1}^{n}\left[\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}\right]\right|.$
@@
$\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}\big|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}\big|+\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon}1.$
@@
$\sup_{(x_{1},x_{2})\in\mathbb{R}^{2}}\frac{1}{n}\sum_{i:\|\tilde{Z}_{i}-\bar{Z}_{i}\|<\varepsilon}\big|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1},\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1},\bar{Z}_{i2}\leq x_{2}\}}\big|$
$\leq\frac{1}{n}\sum_{i:|\tilde{Z}_{i1}-\bar{Z}_{i1}|<\varepsilon}\sup_{x_{1}\in\mathbb{R}}\big|\mathbf{1}_{\{\tilde{Z}_{i1}\leq x_{1}\}}-\mathbf{1}_{\{\bar{Z}_{i1}\leq x_{1}\}}\big|+\frac{1}{n}\sum_{i:|\tilde{Z}_{i2}-\bar{Z}_{i2}|<\varepsilon}\sup_{x_{2}\in\mathbb{R}}\big|\mathbf{1}_{\{\tilde{Z}_{i2}\leq x_{2}\}}-\mathbf{1}_{\{\bar{Z}_{i2}\leq x_{2}\}}\big|.$
Using our assumptions about the true $F$ and because of the convergence indicated by Lemma C.3, this term is $\mathcal{O}_{\mathbb{P},n}(\varepsilon)$.
@@
$\|\tilde{Z}_{i}-\bar{Z}_{i}\|\geq\varepsilon\implies|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\text{ or }|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\varepsilon/\sqrt{2}\}}\leq\gamma.$
@@
$|\tilde{Z}_{i2}-\bar{Z}_{i2}|$ $=\big|\hat{F}_{2\,|\,1,m}^{-1}(u_{i}\,|\,\tilde{Z}_{i1})-\hat{F}_{2\,|\,1,m}^{-1}(v_{i}\,|\,\bar{Z}_{i1})\big|$
$\leq\big|\hat{F}_{2\,|\,1,m}^{-1}(u_{i}\,|\,\tilde{Z}_{i1})-\hat{F}_{2\,|\,1,m}^{-1}(u_{i}\,|\,\bar{Z}_{i1})\big|+\big|\hat{F}_{2\,|\,1,m}^{-1}(u_{i}\,|\,\bar{Z}_{i1})-\hat{F}_{2\,|\,1,m}^{-1}(v_{i}\,|\,\bar{Z}_{i1})\big|.$
@@
$\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\tilde{Z}_{i1}\notin K\}}>\frac{\gamma}{4}\right)\xrightarrow{n\to\infty}0,\quad\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\{\bar{Z}_{i1}\notin K\}}>\frac{\gamma}{4}\right)\xrightarrow{n\to\infty}0.$
@@
$\sup_{\begin{subarray}{c}x_{1},y_{1}\in K\\ |x_{1}-y_{1}|<\delta\end{subarray}}\sup_{u\in[a,b]}\big|\hat{F}_{2\,|\,1,m}^{-1}(u\,|\,x_{1})-\hat{F}_{2\,|\,1,m}^{-1}(u\,|\,y_{1})\big|<\frac{\varepsilon}{2\sqrt{2}}.$
@@
$\sup_{x_{1}\in K}\sup_{\begin{subarray}{c}u,v\in[a,b]\\ |u-v|<\delta_{1}\end{subarray}}\big|\hat{F}_{2\,|\,1,m}^{-1}(u\,|\,x_{1})-\hat{F}_{2\,|\,1,m}^{-1}(v\,|\,x_{1})\big|<\frac{\varepsilon}{2\sqrt{2}}.$
@@
$\left\{|\tilde{Z}_{i2}-\bar{Z}_{i2}|\geq\varepsilon/\sqrt{2}\right\}\subseteq\left\{|\tilde{Z}_{i1}-\bar{Z}_{i1}|\geq\delta\right\}\cup\left\{u_{i}\notin[a,b]\right\}\cup\left\{v_{i}\notin[a,b]\right\}\cup\left\{|u_{i}-v_{i}|\geq\delta_{1}\right\}.$